Solve the equation by completing the square. Round solution to the nearest hundredth if necessary.
X^2 + 6x =2

Found 2 solutions by nerdybill, solver91311:
Answer by nerdybill(7384)

x^2 + 6x =2
take 1/2 of 'b' (6), square it and add it (9) to both sides:
x^2 + 6x + 9 = 2+9
(x+3)(x+3) = 11
(x+3)^2 = 11
x+3 = +-sqrt(11)
x = +-sqrt(11) - 3
x = {-6.317, 0.317}


Answer by solver91311(24713)



You cannot solve



by completing the square. That is because and are two different variables.

It is possible to solve:



So in the hope that is what you meant:

Ordinarily the first step would be to add the opposite of the constant term to both sides, but that has been done already. The second step is to divide thruogh by the lead coefficient. Since the lead coefficient is 1, you can consider that step complete also. Next, divide the coefficient on the first degree term by 2, square the result, and add the result of the squaring operation to both sides of your equation.



Factor the perfect square in the LHS:



Take the square root:



Finish solving:



Just punch the calculator from here.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism